Can you please help me with this problem. We've figured out one answer but there is another answer. So how do you get two answers. Please help us. Rhonda and Tiffany
x^2 + 9x + 18 = 0

Answer by venugopalramana(3286)

x^2 + 9x + 18 = 0
=X^2+3X+6X+18=0
=X(X+3)+6(X+3)=0
=(X+3)(X+6)=0
HENCE X+3=0...THAT IS....X=-3.....OR
X+6=0....THAT IS......X=-6...